# Chirality Amplification and Deracemization in Liquid Crystals

### v1

@@ Abstract
Liquid crystal mesophases of achiral molecules are normally achiral, yet in a few materials they spontaneously deracemize and form right- and left-handed chiral domains. One mechanism that drives deracemization is molecular shape fluctuations between axial chiral conformations, where molecular interactions favor matching chirality and promote helical twist. Cooperative chiral ordering may also play a role in chirality amplification, as when a tiny fraction of chiral dopant drives a nematic phase to become cholesteric. We present a model of cooperative chiral ordering in liquid crystals using Maier-Saupe theory, and predict a phase diagram with a deracemized cholesteric phase as well as racemic nematic and isotropic phases. Our model also predicts chirality amplification in the nematic phase, which may be observed even in materials where the deracemization transition is preempted by a transition to another phase. We compare these results with Monte Carlo simulation studies of the switchable chiral Lebwohl-Lasher model, where each spin switches between right- and left-handed chiral states. Simulation results validate the predicted phase diagram, demonstrate chiral amplification in the racemic nematic phase, and reveal coarsening dynamics in the deracemized phase. Our results suggest that cooperative chiral ordering via molecular shape transitions is a common mechanism in liquid crystals.
@@
Chiral symmetry breaking and chirality amplification are attracting considerable interest in biology, chemistry, materials science and physics due to their relevance to the origin of life and single-handedness of molecules of biological origin [1]. In both, chirality at the molecular scale can be translated to the supramolecular and even the macroscopic scale [2]. For instance, a very small enantiomeric excess of chiral molecular units produces helical polymers of definite handedness, not only in covalent polymers [3] but also in supramolecular ones [4, 5]. Mixing tiny fractions of a chiral compound with achiral molecular building blocks does the same. The former is called the “majority rules” principle and the latter the “sergeants-and-soldiers” effect [6]. Likewise, a small excess of one of the enantiomers or the addition of very small amounts of a chiral compound referred to as “chiral doping” is sufficient to break chiral symmetry in liquid crystals [7, 8]. Recently, lab-synthesized bacteria with biochemical components of opposite chirality to those observed in nature, so-called “mirror cells”, have recently been identified as a potentially dangerous risk to life on Earth [9].
@@
Generally, chirality amplification is thought to arise via co-operative ordering. In polymers, cooperative chiral ordering can occur via long-range correlations between left- and right-handed helically bonded states of neighboring molecular units along the backbone of a polymer chain [4, 10]. These long-range correlations arise because of an underlying Ising-like phase transition that in (quasi) one-dimensional systems is suppressed by thermal fluctuations. We argue here that via a similar mechanism, chiral interactions cause molecular conformations of nearby mesogens in a liquid crystalline fluid to become correlated. This cooperative chiral ordering of molecules gives rise to a diverging chiral susceptibility and must be at the root of chiral symmetry breaking in nematic liquid crystals. We show here that the helical twisting power associated with adding a chiral dopant to a nematic must be inversely proportional to the distance to a second-order phase transition, in which dynamic left-handed and right-handed conformers phase deracemize into cholesteric liquid crystals with opposite handedness. Our calculations show that the helical twisting power may exhibit a complex temperature dependence, dependent on the properties of both the mesogens and the dopants. Our work also explains why some liquid crystalline materials consisting of non-chiral mesogens spontaneously deracemize and form right- and left-handed chiral domains [11], noting also that pretransitional fluctuations signaling deracemization have been seen in simulations of achiral bent-core model mesogens [12]. We are not aware of any experimental work showing that deracemization has yet been observed in more conventional nematic liquid crystals, but argue that the chiral transition that we identify may be hidden below the transition temperature to a smectic or crystalline phase.
@@
We first discuss the main ingredients and findings of our mean-field theory, and next discuss those of our Monte Carlo studies. We end with a short discussion and conclusions, and suggestions for further study.
@@ Results
As a starting point we take the Maier-Saupe theory of van der Meer and collaborators [13], and extend it to mixtures of $\pm$ enantiomers: compounds with axial chiralities P and M that prefer opposite cholesteric wave numbers. The fraction of $\pm$ components we denote $f_{+}=1-f_{-}$, a quantity that we shall treat as an equilibrium variable. We note in this context that, for example, in non-chiral biphenyl mesogens these enantiomers may inter convert dynamically [7] with estimated barriers between 3 and 5 times the thermal energy (at room temperature) [14]. Interactions between the $+$ components favor a dimensionless helical wave number $+q$ and those between the -- components a helical wave number of $-q$. In this model, interactions between $+$ and -- components do not produce a helical pitch. For simplicity we presume all nematic interactions to be the same between the various components and described by a single coupling constant $J>0$. In contrast to the earlier work of Terenteva and Shiyanovskii [15], our model explicitly deals with the chiral interaction between the mesogens and predicts a cholesteric pitch. Our Monte Carlo simulations, discussed in more detail in the next section, are based on the same type of interaction potential [13].
Accounting for the mixing entropy of the two species of compound, we obtain for the Helmholtz free energy $F$ per particle
@@
[…]
The third and fourth terms in Eq. 1 arise due to the chiral interaction that induces a helical deformation of the director field with dimensionless wave number $q$. The quadratic terms in $q$ account for the reduction of the nematic interaction due to the twisting of the director field, whilst the terms linear in $q$ measure how strongly the chiral interaction is favored by the as yet unknown degree of helicity of the director field. These terms follow from a mean-field type of analysis of a chiral Lebwohl-Lasher type of interaction potential between model mesogens [13]; see also Eq. 5.
The last three terms describe the mixing entropy of the species $\pm$, and a term that breaks chiral symmetry with $g$, a free energy difference between the two enantiomers. The factor of 2 we introduce to connect with our simulations in which the equivalent term is $g\eta\propto 2gf_{+}$. This chiral biasing term may, e.g., represent a chemical potential difference that in a grand canonical ensemble regulates the enantiomeric excess $\eta=2f_{+}-1=1-2f_{-}$. Alternatively, it may represent a free energy difference between the two enantiomers, induced by the action of circularly polarized light or the binding of a chiral dopant [7]. In the latter case we expect that $g$ is proportional to the concentration of dopant, the consequences of which we investigate in more detail below.
The optimal orientational distribution function $P(\cos\theta)$ functionally minimizes the free energy, and the equilibrium values of the chiral wave vector $q$ and the fraction $+$ and -- states $f_{+}=1-f_{-}$ also minimize the free energy. For $g=0$ and a sufficiently weak chiral interaction, the transition between a racemic isotropic and a racemic nematic phase occurs at the clearing temperature $T_{\mathrm{IN}}$ for which $J/k_{\mathrm{B}}T_{\mathrm{IN}}=4.54$ where $S=0.43$ in the nematic phase [16], when the free energies of the coexisting isotropic and nematic phase are equal. Let us consider the fluid to be in the nematic phase with temperature $T<T_{\mathrm{IN}}$ and order parameter $S>0.43$.
@@
with $q_{0}=K/2J$ the intrinsic cholesteric wave number of the enantiomers and $x$ the ratio of the equilibrium value of $q$ and that of the intrinsic one, $q_{0}$. We require that $|q_{0}|\lesssim 1$ for the pitch not to drop below a microscopic length, which corresponds to $|K|/J\lesssim 2$. As expected, we find for racemic mixtures, with $f_{+}=f_{-}=1/2$ and enantiomeric excess $\eta=0$, $x=0$ and hence $q=0$. If on the other hand $f_{+}=1-f_{-}=0,1$ and $\eta=\pm 1$, we retrieve $x=\pm 1$ and $q=q_{0}$, which is the equilibrium value of the cholesteric wave number of the enantiomerically pure material within our model.
We also need to minimize the free energy Eq. 1 with respect to $f_{+}=1-f_{-}$, to find
@@
where the strength of the chiral interaction is given by $\chi\equiv(K/J)^{2}S^{2}J/2k_{\mathrm{B}}T=2.27\hskip 2.84544pt(T_{\mathrm{IN}}/T)\hskip 2.84544pt(K/J)^{2}S^{2}$. In the absence of a chiral interaction between the mesogens, $\chi=0$. This may be the case if $K=0$ and $q_{0}=0$, or if the fluid is in the isotropic phase where $S=0$. In both cases, the fractions $+$ and -- enantiomers obey a simple Boltzmann distribution with $f_{+}/f_{-}=\exp(2g/k_{\mathrm{B}}T)$. The corresponding enantiomeric excess $\eta=\tanh g/k_{\mathrm{B}}T$ behaves like the average spin state of the Ising model of ferromagnetism, where $g$ takes on the role of magnetic field. In our model, the spins do not directly couple, only indirectly through the chiral interaction between the species of nematogen.
@@
Figure 1: Left: Enantiomeric excess $\eta$ (blue – inner curve) and ratio $x$ of the cholesteric wave number $q$ and the maximal value $q_{0}$ (red – outer curve) in the co-existing deracemized cholesteric phases as a function of the dimensionless magnitude of the chiral interaction between the nematogens $\chi$ defined in the main text. Right: phase diagram as a function of the temperature $T/T_{\mathrm{IN}}$ scaled to the nematic transition temperature $T_{\mathrm{IN}}$ and the ratio of the chiral and nematic interaction strengths $|K|/J$. The vertical line in red indicates the isotropic-nematic transition. The curved line in blue demarcates the chiral transition between the deracemized cholesteric phase and the racemic nematic phase for $T<T_{\mathrm{IN}}$ and the racemic isotropic phase for $T>T_{\mathrm{IN}}$, which cross at the tricritical point (indicated).
@@
For $g=0$, there is no overall enantiomeric excess: the mesogens have no inherent preference for either enantiomeric state. However, even if $g=0$ local chiral symmetry may still be broken. This happens if $\chi>1$, as is shown in Fig. 1 (left) where we present a numerical solution of Eqs. 2 and 3 for $g=0$. For $\chi\leq 1$ the enantiomeric excess $\eta\equiv 2f_{+}-1$ and cholesteric wave number $x=q/q_{0}$ are both equal to zero. This represents the achiral, racemic nematic phase. For $\chi\geq 1$, local chiral symmetry is broken and the racemic nematic phase deracemizes into two co-existing cholesteric phases with opposite cholesteric wave number ±$\pm|x|$ and opposite enantiomeric excess ±$\pm|\eta|$. In our model, the chiral or deracemization transition is continuous, and is reminiscent of the spontaneous magnetization of the Ising model below the Curie temperature.
@@
The condition $\chi_{\mathrm{c}}=(K/J)^{2}S^{2}(T_{\mathrm{c}})J/2k_{\mathrm{B}}T_{\mathrm{c}}=1$ identifies a critical chiral transition or deracemization temperature $T_{\mathrm{c}}/T_{\mathrm{IN}}=2.27\hskip 2.84544pt(K/J)^{2}S^{2}(T_{\mathrm{c}})$. For weak chiral interaction, so small values of $|K|/J$, $T_{\mathrm{c}}<T_{\mathrm{IN}}$. If $T_{\mathrm{c}}\ll T_{\mathrm{IN}}$ the chiral transition might be obscured by a transition to, say, a smectic or a crystalline phase. On the other hand, $T_{\mathrm{c}}$ moves towards $T_{\mathrm{IN}}$ with increasing chiral interaction $|K|$ as is shown in our phase diagram Fig. 1 (right). The deracemization transition crosses the nematic transition at a tricritical point for $|K|/J=1.54$. For $|K|/J>1.54$ there is a direct transition from a racemic isotropic phase to two deracemized cholesteric phases of equal volume. The phase diagram we obtain from our simulations, shown in Fig. 3 (top), is in good agreement with that from our mean-field theory of Fig. 1 (right).
@@
Not surprisingly, we obtain mean-field critical exponents near the chiral transition. In fact, for $T\rightarrow T_{\mathrm{c}}$, Eqs. 2 and 3 can be solved analytically giving $x\sim\pm 4\hskip 2.84544pt(3/28)^{1/2}(\chi-1)^{1/2}\pm\propto(T_{\mathrm{c}}-T)^{1/2}$ and $\eta\sim\pm 2\hskip 2.84544pt(3/28)^{1/2}(\chi-1)^{1/2}\propto\pm(T_{\mathrm{c}}-T)^{1/2}$ for $\chi\geq 1$ or $T\leq T_{\mathrm{c}}$, and $\eta=x=0$ for $\chi\leq 1$ or $T>T_{\mathrm{c}}$. Arguably, the existence of a second order chiral transition explains the phenomenon of chirality amplification because near a critical point susceptibilities tend to become large and in fact diverge at the critical point. This should show up in the response of the chiral order parameters $x$ and $\eta$ to the chiral biasing field $g$.
@@
This point is illustrated in Fig. 2 (a). The figure shows our solution to Eqs. 2 and 3 for the enantiomeric excess $\eta$ as a function of $g/k_{\mathrm{B}}T$ for different scaled temperatures $T/T_{\mathrm{IN}}$ and fixed chiral interaction $K/J=0.333$. Recall that the chiral interaction strength $\chi$ can be written as $\chi=\frac{1}{2}(K/J)^{2}S^{2}(J/k_{\mathrm{B}}T)$ where $J/k_{\mathrm{B}}T=4.54\hskip 2.84544ptT_{\mathrm{IN}}/T$. This means that the enantiomeric excess $\eta$ is not only a function of the dimensionless biasing field $g/k_{\mathrm{B}}T$, but also a function of $K/J$ as well as of $T/T_{\mathrm{IN}}$ because of the nematic order parameter $S=S(T/T_{\mathrm{IN}})$ also depends on the temperature. The values of $S(T/T_{\mathrm{IN}})$ we obtain from a numerical solution of the self-consistent field equation of Maier-Saupe theory. For the temperatures $T/T_{\mathrm{IN}}$ shown, $\chi$ is smaller than unity and the nematic remains racemic. Still, we do see a significant increase in the enantiomeric excess relative to the hyperbolic tangent relation describing the case $K=0$, and more so the lower the temperature. The trends of our theoretical predictions agree with those from our simulations, as Fig. 2 (b) confirms. See also the Supporting Information, Fig. S1, in which we investigate the impact of larger and smaller values of $K/J$. For $\chi>1$, the Maier-Saupe theory predicts hysteresis of the enantiomeric excess $\eta$ as a function of the biasing field $g$, with metastable positive (negative) valued $\eta$ for negative (positive) values of $g$ (not shown).
@@
In the limit $\chi\rightarrow 1$ the susceptibilities of the enantiomeric excess $\chi_{\eta}\equiv\lim_{g\rightarrow 0}\partial\eta/\partial g$ and that of the cholesteric wave number $\chi_{x}\equiv\lim_{g\rightarrow 0}\partial x/\partial g$ both actually diverge. From Eqs. 2 and 3, we find that for $\chi\uparrow 1$, $\chi_{\eta}=\frac{1}{2}\chi_{x}=\frac{1}{4}(1-\chi)^{-1}/k_{\mathrm{B}}T$, whilst for $\chi\downarrow 1$, $\chi_{\eta}=\frac{1}{2}\chi_{x}\sim\frac{7}{24}(\chi-1)^{-1}/k_{\mathrm{B}}T$. So, in both cases we find $\chi_{\eta}\propto|T-T_{\mathrm{c}}|^{-1}$. Our simulations produce the same mean-field exponent for $\chi_{\eta}$ (Fig. S4) albeit that we have not performed any finite-size analysis to obtain an accurate value. Finally, accounting in the free energy for square gradient terms in the variables $x$ and $f_{+}=(\eta+1)/2$ produces two correlation lengths proportional to $1/\sqrt{1-\chi}$ confirming that chiral conformational states are indeed “contagious” and passed on to neighboring mesogens through the chiral interaction.
The susceptibility $\chi_{x}$ associated with the cholesteric wave number may be interpreted to represent what is known as the helical twisting power of a chiral dopant [17]. A tiny induced preference for a particular enantiomeric form of the nematogens becomes strongly enhanced by the coupling to the chiral symmetry breaking transition. This may well explain why so little chiral dopant is needed to turn a nematic into a cholesteric [7, 8]. That different dopants produce varying helical twisting powers if added to the same mesogen may be understood in terms of unequal surface interactions with the director field [18], expressed in our model in different values of the binding free energy, $g$. In fact, for large enough binding energies, $g$ may be replaced by $gX$ with $X\ll 1$ the mole fraction of dopant, in which case the helical twisting power, $\beta$, may be written as

### v2

# Chirality Amplification and Chiral Segregation in Liquid Crystals

@@ Abstract
Liquid crystal mesophases of achiral molecules are normally achiral, yet in a few materials they spontaneously segregate and form right- and left-handed chiral domains. One mechanism that drives chiral segregation is molecular shape fluctuations between axial chiral conformations, where molecular interactions favor matching chirality and promote helical twist. Cooperative chiral ordering may also play a role in chirality amplification, as when a tiny fraction of chiral dopant drives a nematic phase to become cholesteric. We present a model of cooperative chiral ordering in liquid crystals using Maier-Saupe theory, and predict a phase diagram with a segregated cholesteric phase with alternating domains of left- and right-handed chiral twist, with opposite enantiomeric excess, in addition to racemic nematic and isotropic phases. Our model also demonstrates how chiral molecular fluctuations influence the helical twisting power of dopants in the nematic phase, which may be observed even in materials where the segregated cholesteric phase is preempted by a transition to another phase. We compare these results with Monte Carlo simulation studies of the switchable chiral Lebwohl-Lasher model, where each spin switches between right- and left-handed chiral states. Simulation results validate the predicted phase diagram, demonstrate chiral amplification in the racemic nematic phase, and reveal complex coarsening dynamics in the segregated cholesteric phase. These results suggest that molecular fluctuations between degenerate chiral configurations may be a common mechanism to produce cooperative chiral order in achiral liquid crystals.
@@
Chiral symmetry breaking and chirality amplification are attracting considerable interest in biology, chemistry, materials science and physics due to their relevance to the origin of life and single-handedness of molecules of biological origin [4]. In both, chirality at the molecular scale can be translated to the supramolecular and even the macroscopic scale [45]. For instance, a very small enantiomeric excess of chiral molecular building blocks produces helical polymers of definite handedness, not only in covalent polymers [23] but also in supramolecular ones [49, 25]. Mixing tiny fractions of a chiral compound with achiral molecular building blocks does the same [31, 22]. Likewise, a small excess of one of the enantiomers or the addition of very small amounts of a chiral compound referred to as “chiral doping” is sufficient to break chiral symmetry in liquid crystals [13, 41]. Recently, lab-synthesized bacteria with biochemical components of opposite chirality to those observed in nature, so-called “mirror cells”, have been identified as a potential risk to life on Earth [1].
@@
Generally, chirality amplification is thought to arise via cooperative ordering. In polymers, cooperative chiral ordering can occur via long-range correlations between left- and right-handed helically bonded states of neighboring molecular units along the backbone of a polymer chain [49, 38]. These long-range correlations arise because of an underlying Ising-like phase transition that in (quasi) one-dimensional systems is suppressed by thermal fluctuations. We argue here that via a similar mechanism, chiral interactions cause molecular conformations of nearby mesogens in a liquid crystalline fluid to become correlated or “synchronized” [44]. This cooperative chiral ordering of molecules gives rise to a diverging chiral susceptibility and must be at the root of chiral symmetry breaking in nematic liquid crystals. We show here that the helical twisting power associated with adding a chiral dopant to a nematic must be inversely proportional to the distance to a second-order phase transition, in which dynamic left-handed and right-handed conformers phase separate into cholesteric liquid crystals with opposite handedness. Our calculations show that the helical twisting power may exhibit a complex temperature dependence.
Our work also explains how and why a wide variety of liquid crystalline materials consisting of achiral mesogens can exhibit chiral segregation, also known as spontaneous resolution, to form coexisting right- and left-handed chiral domains [33, 3]. Although spontaneous resolution has not been seen in “conventional” nematic liquid crystals such as $n$-CB or MBBA, with the exception of, e.g., alkyl benzoic acid compounds [43], we put forward that for such mesogens it may be obscured by a transition to a smectic or a crystalline phase. Even if obscured by a transition to another phase, pre-transitional effects should still be noticeable in the helical twisting power. Of particular interest for the present work is chiral segregation in fluids of achiral mesogenic compounds that exhibit a so-called nematic twist-bend phase [45, 30]. Like the cholesteric phase, this phase characterized by a helical modulation of the director field but with very much smaller pitch. Our theory provides an explanation for why such compounds are susceptible to chiral segregation.
@@
We first discuss the main ingredients and findings of our mean-field theory, and next discuss those of our Monte Carlo studies. We end with a discussion and conclusions, and suggestions for further study.
Methods
Theory
Following van der Meer et al[46], we find from a mean-field analysis of the impact of a chiral Lebwohl-Lasher type of interaction potential between model mesogens to lead to expressions for the self-consistent fields experienced by species $\alpha=\pm$ due to the presence of species $\beta=\pm$ of the form
The chiral coupling constants $K_{\alpha,\beta}$ are $K_{+,+}=K$ between the ++ species, $K_{-,-}=-K$ between the - species and $K_{+,-}=K_{-,+}=0$ between the species with opposite handedness. In Eq. 1, we have absorbed unimportant numerical constants in the definitions of the interaction strengths $J$ and $K$ as well as the cholesteric wave number $q$ that in our description is made dimensionless by multiplication with a microscopic length scale [46]. The free energy of interaction follows from summing all self-consistent fields experienced by species $\alpha$ multiplied with the appropriate fraction $f_{\alpha}$, and accounting for double counting. Adding the Gibbs entropy for the orientational distribution and that for the distribution over the species $\alpha$, as well as the contribution of the biasing field $g$, then produces, after some rearranging, Eq. 2.
Simulation Methods
To explore cooperative chiral ordering and spontaneous chiral segregation by means of Monte Carlo computer simulation, we study a switchable chiral 3D Lebwohl-Lasher model based on the Hamiltonian of Eq. 6. We perform Monte Carlo simulations of the model via the Metropolis algorithm, where at each Monte Carlo step we separately vary the local orientation $\hat{u}_{i}$ and chiral state parameter $\eta_{i}$. We use a novel method for trial spin reorientation which realizes a smoother reorientation of the spin, rather than the Barker-Watts technique [2] used in [40]. To perform a random walk on a unit sphere, we calculate two vectors orthogonal to the orientation vector $\hat{u}_{i}$ and tangent to the unit sphere, then reorient the spin to a random position on the edge of a disk of radius $d$, and normalize the orientation vector to unit length. The disk size $d$ represents the step size and is adjusted to keep the Monte Carlo acceptance rate in the range of 0.4 to 0.6 during equilibration at each temperature, and after that is held fixed. We initialize each simulation in the racemic isotropic phase.
Lattice-based Monte-Carlo simulations are a computationally “embarrassingly parallelizable” problem.Our Monte Carlo code is highly optimized to run on graphics processing units (GPUs) using a combination of Fortran and OpenACC for the bulk of the computation, and Python for simulation management, data processing, analysis, and visualization. Simulations were performed on the National Center for Supercomputing Applications’ Delta cluster through the NSF ACCESS program [5]. A copy of the simulation code is available [8].
@@ Results
Our starting point is the Maier-Saupe theory of van der Meer and collaborators [46] that we extend to mixtures of $\pm$ enantiomers: compounds with axial chiralities P and M, and opposite cholesteric pitches. The fraction of $\pm$ components $f_{+}=1-f_{-}$ we treat as an equilibrium variable. For example, in non-chiral biphenyl mesogens such as, say, 5CB these stereo isomers may inter convert dynamically [13] with estimated barriers a few times the thermal energy [52]. Interactions between equal pairs of $\pm$ components favor a dimensionless helical wave number $\pm q$. In this model, interactions between ++ and - components do not produce a helical pitch, and hetero-chiral interactions are less strong than homo-chiral ones. For simplicity all nematic interactions between the various components are equal, and described by a single coupling constant $J>0$. In contrast to the earlier work [42], our model explicitly deals with chiral interactions between the mesogens and predicts a cholesteric pitch. Our Monte Carlo simulations, discussed in more detail in the next section, are based on the same type of interaction potential [46].
Accounting for the mixing entropy of the two species of compound, we obtain for the Helmholtz free energy $F$ per mesogen
@@
where $k_{\mathrm{B}}T$ denotes the thermal energy with $k_{\mathrm{B}}$ Boltzmann’s constant and $T$ the absolute temperature. The first two terms represent the classical expression for the Maier-Saupe free energy (per molecule) describing the isotropic-nematic phase transition [48]. The first describes the Gibbs entropy associated with the distribution $P=P(\cos\theta)$ over the polar angle $\theta$ between a mesogen and the local director, and the second the contribution from the nematic interaction between the mesogens. In the latter, $S=\langle P_{2}(\cos\theta)\rangle\in[-1/2,+1]$ denotes the scalar nematic order parameter taken to be the same for both $\pm$ components, with $P_{2}(\cos\theta)=\frac{3}{2}(\cos\theta)^{2}-\frac{1}{2}$ the second Legendre polynomial. The angular brackets $\langle\cdots\rangle$ indicate an orientational average $\int_{-1}^{+1}d\cos\theta\hskip 2.84544ptP(\cos\theta)(\cdots)$ over the orientational distribution function $P(\cos\theta)$ of the nematogens. We ignore any biaxility in the distribution function that may arise in the cholesteric phase.
The third and fourth terms in Eq. 2 arise due to the chiral interaction that induces a helical deformation of the director field with dimensionless wave number $q$. The quadratic terms in $q$ account for the reduction of the nematic interaction in the twisted director field, while the terms linear in $q$ measure how strongly the chiral interaction is enhanced. These terms follow from a mean-field analysis of a chiral Lebwohl-Lasher type of interaction potential between model mesogens [46]; see also Eq. 6.
The last three terms describe the mixing entropy of the species $\pm$, and a term that breaks chiral symmetry with $g$ a free energy difference between the two enantiomers. This chiral biasing term may, e.g., represent a chemical potential difference that in a grand canonical ensemble regulates the enantiomeric excess $\eta=2f_{+}-1=1-2f_{-}$. Alternatively, it may be a free energy difference between the two enantiomers induced by the action of circularly polarized light or the interaction with a chiral dopant [13]. In the latter case we expect that $g$ is proportional to the concentration of dopant, the consequences of which we investigate in more detail below. See the SI for a discussion.
The optimal orientational distribution function $P(\cos\theta)$ minimizes the free energy, as do the equilibrium values of the chiral wave vector $q$ and the fraction ++ and - states $f_{+}=1-f_{-}$. For $g=0$ and a sufficiently weak chiral interaction, the transition between a racemic isotropic and a racemic nematic phase occurs at the clearing temperature $T_{\mathrm{IN}}$ for which $J/k_{\mathrm{B}}T_{\mathrm{IN}}=4.54$ where $S=0.43$ in the nematic phase [48], when the free energies of the coexisting isotropic and nematic phase are equal. Let us consider the fluid to be in the nematic phase, with temperature $T<T_{\mathrm{IN}}$ and order parameter $S>0.43$.
@@
with $q_{0}=K/2J$ the intrinsic cholesteric wave number of the enantiomers and $x$ the ratio of the equilibrium value of $q$ and $q_{0}$. We require that $|q_{0}|\lesssim 1$ or $|K|/J\lesssim 2$ for the pitch not to drop below a microscopic length.
We also need to minimize the free energy Eq. 2 with respect to $f_{+}=1-f_{-}$, and find
@@
where the strength of the chiral interaction is given by $\chi\equiv(K/J)^{2}S^{2}J/2k_{\mathrm{B}}T=9.08\hskip 2.84544ptq_{0}^{2}\hskip 2.84544ptS^{2}\hskip 2.84544pt(T_{\mathrm{IN}}/T)$. The temperature dependence of $S=S(T/T_{\mathrm{IN}})$ follows from the self-consistent field equation of Maier-Saupe theory [48]. In the absence of a chiral interaction between the mesogens, $\chi=0$. This happens if $K=0$ and $q_{0}=0$, or if the fluid is in the isotropic phase and $S=0$. The fractions $\pm$ enantiomers then obey a simple Boltzmann distribution, giving for the enantiomeric excess $\eta=\tanh g/k_{\mathrm{B}}T$, not unlike the average spin state of the Ising model of ferromagnetism where $g$ takes on the role of magnetic field. In our model, the “spins” do not directly couple, only indirectly – see Eq. 6.
@@
Figure 1: Left: Enantiomeric excess $\eta$ (blue – inner curve) and ratio $x$ of the cholesteric wave number $q$ and the maximal value $q_{0}$ (red – outer curve) in the co-existing segregated cholesteric phases as a function of the dimensionless magnitude of the chiral interaction between the nematogens $\chi$. Right: phase diagram as a function of the temperature $T/T_{\mathrm{IN}}$ scaled to the nematic transition temperature $T_{\mathrm{IN}}$, and the ratio of the chiral and nematic interaction strengths $|K|/J$. The vertical line in red indicates the isotropic-nematic transition. The curved line in blue demarcates the chiral transition between the segregated cholesteric phase and the racemic nematic phase for $T<T_{\mathrm{IN}}$ and the racemic isotropic phase for $T>T_{\mathrm{IN}}$, which cross at what we interpret to be a critical end point (CEP) indicated by the circle. See also the main text.
@@
For $g=0$, there cannot be an overall enantiomeric excess as the mesogens have no preference for either enantiomeric state. However, local chiral symmetry may still be broken. This happens if $\chi>1$, as is shown in Fig. 1 (left), having numerically solved Eqs. 3 and 4 for $g=0$. For $\chi\leq 1$ the enantiomeric excess $\eta\equiv 2f_{+}-1$ and cholesteric wave number $x=q/q_{0}$ are both equal to zero. This represents the achiral, racemic nematic phase. For $\chi\geq 1$, local chiral symmetry is broken and the racemic nematic phase segregates into two co-existing cholesteric phases with opposite cholesteric wave number $\pm|x|$ and opposite enantiomeric excess $\pm|\eta|$. In our model, the chiral or mirror-image symmetry breaking transition is continuous and is reminiscent of the spontaneous magnetization of a ferromagnet below the Curie temperature.
@@
The condition $\chi_{\mathrm{c}}=9.08\hskip 2.84544ptq_{0}^{2}\hskip 2.84544ptS^{2}_{\mathrm{c}}\hskip 2.84544ptT_{\mathrm{IN}}/T_{\mathrm{c}}=1$ identifies a critical chiral transition or chiral segregation temperature $T_{\mathrm{c}}/T_{\mathrm{IN}}=9.08\hskip 2.84544ptq_{0}^{2}S^{2}_{\mathrm{c}}$ with $S_{\mathrm{c}}$ the nematic order parameter at $T=T_{\mathrm{c}}$. For weak chiral interaction, implying small values of $q_{0}=|K|/2J$, $T_{\mathrm{c}}\ll T_{\mathrm{IN}}$. In that case the chiral transition might be obscured by a transition to a smectic or a crystalline phase. On the other hand, $T_{\mathrm{c}}$ moves towards $T_{\mathrm{IN}}$ with increasing chiral interaction strength, as is shown in Fig. 1 (right). The chiral transition crosses the nematic transition at for $|K|/J=1.54$. For $|K|/J>1.54$ there is a direct transition from a racemic isotropic phase to two segregated cholesteric phases of opposite handedness and enantiomeric excess, and equal volume. This arguably marks a critical end point. (See also below.) The phase diagram we obtain from our simulations, Fig. 3 (top), is in good agreement with that from our mean-field theory, Fig. 1 (right).
@@
Not surprisingly, we obtain mean-field critical exponents near the chiral transition. For $T\rightarrow T_{\mathrm{c}}$, Eqs. 3 and 4 can be solved exactly, giving $x\sim\pm 4\hskip 2.84544pt(3/28)^{1/2}(\chi-1)^{1/2}\pm\propto(T_{\mathrm{c}}-T)^{1/2}$ and $\eta\sim\pm 2\hskip 2.84544pt(3/28)^{1/2}(\chi-1)^{1/2}\propto\pm(T_{\mathrm{c}}-T)^{1/2}$ for $\chi\geq 1$ or $T\leq T_{\mathrm{c}}$, and $\eta=x=0$ for $\chi\leq 1$ or $T>T_{\mathrm{c}}$. Arguably, the existence of a second order chiral transition explains the phenomenon of chirality amplification because thermodynamic susceptibilities tend to become large and even diverge upon approach of the critical point. This shows up in the response of the chiral order parameters $x$ and $\eta$ to the chiral biasing field $g$.
This point is illustrated in Fig. 2 (a). The figure shows our solution to Eqs. 3 and 4 for the enantiomeric excess $\eta$ as a function of $g/k_{\mathrm{B}}T$ for different scaled temperatures $T/T_{\mathrm{IN}}$ and fixed chiral interaction $K/J=0.333$. Clearly, the enantiomeric excess $\eta$ turns out not only a function of the dimensionless biasing field $g/k_{\mathrm{B}}T$ but also of $T/T_{\mathrm{IN}}$. The reason is that the chiral interaction strength $\chi$ is a function not only of $K/J$ but also of $T/T_{\mathrm{IN}}$.
@@
For the temperatures $T/T_{\mathrm{IN}}$ shown, $\chi$ is smaller than unity and the fluid remains in the racemic nematic phase. Still, we do see a significant increase in the enantiomeric excess relative to the hyperbolic tangent relation describing the case $K=0$, and more so the lower the temperature. The trends of our theoretical predictions agree with those from our simulations, as Fig. 2 (b) confirms. See also the Supporting Information, Fig. S1, in which we investigate the impact of larger and smaller values of $K/J$. For $\chi>1$, the Maier-Saupe theory predicts hysteresis of the enantiomeric excess $\eta$ as a function of the biasing field $g$, with metastable positive (negative) valued $\eta$ for negative (positive) values of $g$ (shown in SI).
@@
In the limit $\chi\rightarrow 1$ the susceptibilities of the enantiomeric excess $\chi_{\eta}\equiv\lim_{g\rightarrow 0}\partial\eta/\partial g$ and of the cholesteric wave number $\chi_{x}\equiv\lim_{g\rightarrow 0}\partial x/\partial g$ both diverge. From Eqs. 3 and 4, we find that for $\chi\uparrow 1$, $\chi_{\eta}=\frac{1}{2}\chi_{x}=\frac{1}{4}(1-\chi)^{-1}/k_{\mathrm{B}}T$, while for $\chi\downarrow 1$, $\chi_{\eta}=\frac{1}{2}\chi_{x}\sim\frac{7}{24}(\chi-1)^{-1}/k_{\mathrm{B}}T$. So, in both cases we find $\chi_{\eta}\propto|T-T_{\mathrm{c}}|^{-1}$. Our simulations produce approximately the same mean-field exponent (Fig. S4) although we have not performed a finite-size analysis to obtain a more accurate value. Finally, accounting in the free energy for square gradient terms in the variables $x$ and $f_{+}=(\eta+1)/2$ produces two correlation lengths proportional to $1/\sqrt{1-\chi}$ confirming that chiral conformational states are indeed in a way contagious and transferred to nearby mesogens through the chiral interaction.
The susceptibility $\chi_{x}$ associated with the cholesteric wave number may be interpreted to represent what is known as the helical twisting power of a chiral dopant [17]. A tiny induced preference for a particular enantiomeric form of the nematogens becomes strongly enhanced by the coupling to the chiral symmetry breaking transition. This may well explain why so little chiral dopant is needed to turn a nematic into a cholesteric [13, 41]. Differences in the helical twisting power associated with different dopants may be understood in terms of unequal surface interactions with the director field [18], expressed in our model in different values of the interaction free energy, $g$. In fact, as we show in the SI for large enough binding free energies, $g$ may be replaced by $gX$ with $X\ll 1$ the mole fraction of dopant, in which case the helical twisting power, $\beta$, can be written as